James Bond was chasing an evil villain to the top of the crane 100 point meters above the ground the villain leaps off and lands on the rooftop that is 9.0 m away horizontally the rooftop below is 65 m above the ground if bond leaps horizontally with a speed of 3.5 m/s will 007 come up short and plummet to his impaling doom or will he live another day
1
Expert's answer
2021-03-26T11:39:04-0400

Find the time it will take to fall from the height of 100 m to 65 m:


"t=\sqrt{\frac{2y}{g}}=\sqrt{\frac{2(H-h)}{g}}."

Find the distance the 007 can go along the horizontal during the time t:


"d=vt=v\sqrt{\frac{2(H-h)}{g}}=9.35\text{ m}."

Greater than 9 m, so, he must be safe and sound.